OpenStudy (anonymous):

solve. x- (3square root of x) -4=0

jhonyy9 (jhonyy9):

x-3sqrtx -4 =0 x-4=3sqrtx (x-4)^2 =9x x^2 -8x +16=9x x^2 -17x +16=0 (x-16)(x-1)=0 x-16=0 x=16 x-1=0 x=1

OpenStudy (anonymous):

nice!